# Low-Latency and Low-Complexity MLSE for Short-Reach Optical Interconnects

### v1

@@ IV Simplified L2S-MLSE
Since the latency of L2S-MLSE is reduced at the cost of complexity, simplified L2S-MLSE is proposed in this section. Simplified L2S-MLSE consists of two parts: computational simplification and reduced state. Computational simplification is proposed to reduce the number of multipliers. Reduced state is combined with computational simplification, further reducing the number of adders and comparators.
@@ IV-A Computational Simplification
1) BM Calculation and ACS of the First Layer: To reduce the number of multipliers, the computational simplification is proposed, which combines the common terms in the BM calculation and addition operations of the first layer. The 64 expansion equations of the BM calculation and addition operations from time index n1n-1 to time index $n+1$ in the first layer are shown in Appendix A. These 64 expansion equations can be expressed as the sum of vectors $\bf{A}$}, $\bf{B}$}, $\bf{C}$}, and $\bf{D}$}. Table A.I in Appendix A provides the detailed definitions of these vectors. Both the vector ${\bf{A}}=[{A_{1}},{A_{2}},...,{A_{16}}]$ and the vector ${\bf{C}}=[{C_{1}},{C_{2}},...,{C_{16}}]$ are related to $\alpha$ and constant values. Both the vector ${\bf{B}}=[{B_{1}},{B_{2}},...,{B_{16}}]$ and the vector ${\bf{D}}=[{D_{1}},{D_{2}},...,{D_{16}}]$ are related to the received symbol $y$.
Figure 5: Illustration of 2-step trellis from time index n1n-1 to time index $n+1$ in the first layer of L2S-MLSE.
@@
Through combining the common terms, the vectors $\bf{A}$} and $\bf{C}$} include 1 variable multiplier for $\alpha^{2}$, 4 constant multipliers for $10\alpha^{2}$, $18\alpha^{2}$, $12\alpha$, $36\alpha$, and additional 31 adders. Multiplication with powers of 2 can be achieved through bit-shift operations, thereby avoiding the use of multipliers. For the calculation of $\bf{B}$} and $\bf{D}$}, $\bf{B}$} can be decomposed into ${\bf{M}}+{\bf{N}}+{\bf{J}}$, while $\bf{D}$} can be decomposed into ${\bf{M}}+{\bf{N}}+{\bf{K}}$. $\bf{M}$} corresponds to the common term $y_{n}+\alpha y_{n+1}$ and its multiples. $\bf{N}$} corresponds to the common term $\alpha y_{n}$ and its multiples. $\bf{J}$} is equal to the common term $6y_{n+1}$ for $\bf{B}$}, and $\bf{K}$} is equal to the common term $2y_{n+1}$ for $\bf{D}$}. Then, the calculation of $\bf{B}$} and $\bf{D}$} can be represented as 5 steps:
  • Computing 2 variable multipliers: $\alpha y_{n}$ and $\alpha y_{n+1}$;
  • Computing 1 adder: common term $y_{n}+\alpha y_{n+1}$;
  • Computing 3 constant multipliers: $6y_{n}+6\alpha y_{n+1}$, $6\alpha y_{n}$ and $6y_{n+1}$;
  • Computing 4 adders for ${\bf{N}}+{\bf{J}}$ and 4 adders for ${\bf{N}}+{\bf{K}}$;
  • Computing 16 adders for ${\bf{M}}+{\bf{N}}+{\bf{J}}$ and 16 adders for ${\bf{M}}+{\bf{N}}+{\bf{K}}$.
@@
Therefore, through combining the common terms, the vectors $\bf{B}$} and $\bf{D}$} include 2 variable multipliers, 3 constant multipliers, and 41 adders. The computation of $\bf{B}$} and $\bf{D}$} can be performed simultaneously with that of $\bf{A}$} and $\bf{C}$}. In ACS, 64 adders are required from Eq. (A.1) to Eq. (A.16) shown in Appendix A, while 48 comparators are utilized after the addition operations. In conclusion, with computational simplification, the overall complexity of BM calculation and ACS operation in the first layer includes $N+1$ variable multipliers, $3N/2+4$ constant multipliers, $105N/2+31$ adders, and $24N$ comparators.
@@
2) ACS of the Other Layers and Survivor Path Selection: The difference between L2S-MLSE with and without computational simplification only exists in the first layer, since the second layer can only utilize the computationally simplified results from the first layer. From the second layer to the last layer, the computing process of L2S-MLSE with computational simplification is identical to that of L2S-MLSE without computational simplification, which preserves the same latency and computational complexity.
@@
To make a fair comparison, the computational simplification is also applied to 1S-MLSE. The expansion equations of the BM calculation from time index n1n-1 to time index $n$ are shown in Appendix B. The expansion equations can be expressed as the sum of vectors $\bf{E}$} and $\bf{F}$}. Table B.I in Appendix B provides the detailed definitions of these vectors. By combining common terms, the vector ${\bf{E}}=[{E_{1}},{E_{2}},...,{E_{8}}]$ includes 1 variable multiplier for $\alpha^{2}$, 3 constant multipliers for $9\alpha^{2}$, $6\alpha$, $18\alpha$, and 12 adders. The vector ${\bf{F}}=[{F_{1}},{F_{2}},...,{F_{8}}]$ includes 1 variable multiplier for $\alpha y_{n}$, 2 constant multipliers for $6\alpha y_{n}$ and $6y_{n}$, plus 8 adders. 16 adders are needed for the addition between $\bf{E}$} and $\bf{F}$}, while 12 comparators are utilized after the addition operations. After the BM calculation with computational simplification, ACS and survivor path selection are the same as those of 1S-MLSE without computational simplification, which preserves the same latency and computational complexity.
Figure 6: BER performance of 1S-MLSE, L2S-MLSE, and their counterparts with computational simplification (CS).
@@
In conclusion, the overall computational complexity of 1S-MLSE with computational simplification is $N+1$ variable multipliers, $2N+3$ constant multipliers, 40N440N-4 adders, and $12N+3$ comparators, while that of L2S-MLSE with computational simplification is $N+1$ variable multipliers, $3N/2+4$ constant multipliers, 169N/233169N/2-33 adders, and 48N3348N-33 comparators. Both 1S-MLSE with computational simplification and L2S-MLSE with computational simplification use the same number of variable multipliers. Because constant multipliers and comparators can be implemented by adders, variable multipliers occupy far more hardware resources than constant multipliers, adders, or comparators. When using computational simplification, the number of variable multipliers is reduced from $O(NM^{L})$ to $O(N)$. The BER performance of 1S-MLSE, 1S-MLSE with computational simplification, L2S-MLSE, and L2S-MLSE with computational simplification is shown in Fig. 6. All four algorithms have the same BER performance. By combining common items, the calculation accuracy remains unchanged, ensuring the computational simplification does not degrade the BER performance.
@@
IV-B Simplified L2S-MLSE with the Combination of Computational Simplification and Reduced State
Figure 7: Block diagram of DSP flows with reduced-state MLSE. The pre-decision is from FFE and the reduced states are used for BM calculation.
@@ IV-B Simplified L2S-MLSE with the Combination of Computational Simplification and Reduced State
To reduce the number of adders and comparators, the reduced state is combined with computational simplification. In our paper, we refer to L2S-MLSE with the combination of computational simplification and reduced state as simplified L2S-MLSE. As shown in Fig. 7, the output signal $d$ of FFE can be regarded as the pre-decision value of the received signal [hoon]. For PAM4 with 4 states of $\{-3,-1,1,3\}$, if the number of states is reduced from 4 to 3 through pre-decision value, the possible states include $\{-3,-1,1\}$ or $\{-1,1,3\}$, which are represented by $\{a,b,c\}$ in Fig. 8(a). If the number of states is reduced from 4 to 2 through pre-decision value, the possible states include $\{-3,-1\}$, $\{-1,1\}$ or $\{1,3\}$, which are represented by $\{a,b\}$ in Fig. 8(b). As shown in Fig. 8(a), when the state number is reduced to 3, the number of ACS units in a 2-step calculation is reduced from 16 to 9, in which only 27 adders and 18 comparators are needed. As shown in Fig. 8(b), when the state number is reduced to 2, the number of ACS units in a 2-step calculation is reduced from 16 to 4, in which only 8 adders and 4 comparators are needed.
@@
Figure 8: Illustration of 2-step trellis in the first layer of L2S-MLSE when the state number is reduced from (a) 4 to 3 or (b) 4 to 2.
2) ACS of the Other Layers and Survivor Path Selection: From the second layer to the last layer, due to the decrease in ACS units, 27 adders and 18 comparators are needed for a 2-step calculation with 3 states. 8 adders and 4 comparators are needed for a 2-step calculation with 2 states. In the final survivor path selection step, the path with the minimal accumulated metrics is chosen from 9 paths for the state number of 3, reducing the number of comparators to 8. Similarly, the path with the minimal accumulated metrics is chosen from 4 paths for the state number of 2, reducing the number of comparators to 3.
Figure 9 compares the BER performance of L2S-MLSE under three configurations: all 4 states (with only computational simplification), simplified L2S-MLSE with 3 states, and simplified L2S-MLSE with 2 states. In our experiment, the noise enhancement caused by FFE is not severe enough to damage the pre-decision value from FFE for 3 or 2 reduced states. Therefore, the BER performance remains nearly identical regardless of whether all 4 states, 3 states, or 2 states are used. For the system with a larger $\alpha$ value, due to the larger noise in the pre-decision value, the BER performance of 2 states may have worse BER performance than 3 states [hoon]. Owing to the nearly identical BER performance in Fig. 9, the simplified scheme in our paper refers to the combination of computational simplification and 2 reduced states. To make a fair comparison, the simplified scheme is also applied to 1S-MLSE. The overall computational complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE is summarized in Table I.
TABLE I: Computational complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE
1S-MLSE Simplified 1S-MLSE L2S-MLSE Simplified L2S-MLSE
Variable Multiplier $16N$ $N+1$ $16N$ $N+1$
Constant Multiplier 11 $2N+3$ 11 $3N/2+4$
Adder $32N$ $16N+8$ 80N4880N-48 $57N/2+23$
Comparator $12N+3$ $2N+1$ 48N3348N-33 4N14N-1
Figure 9: BER performance of L2S-MLSE with computational simplification and different numbers of reduced states.
@@ V Experimental Setup
The experiment of a 112-Gbit/s PAM4 signal over 2-km SSMF transmission in C-band is performed to verify the performance of the proposed MLSE. The experimental setup is shown in Fig. 10. At the transmitter, $4\times 10^{5}$ pseudo-random binary sequences (PRBS) are mapped to $2\times 10^{5}$ PAM4 symbols, in which the first 1000 symbols are served as training symbols. To improve the transmission performance, the optimized pre-equalization method proposed in our previous work is applied to compensate for partial bandwidth-limited distortion [guo]. The PAM4 signal is pulse-shaped by a root-raised cosine (RRC) filter with the roll-off factor of 0.1. Then, the signal is resampled by 1.25 sps to align with the sampling rate of the digital-to-analog converter (DAC). The 56-GBaud PAM4 analog signal is generated by a DAC with a sampling rate of 70 GSa/s. The 3-dB bandwidth of the 8-bit DAC is 16 GHz. Afterwards, the analog signal is amplified by a 40-Gbps electrical amplifier (EA, CENTELLAX OA4MVM3). A 40-Gbps optical electro-absorption modulator (EAM, OM5757C-CTM388) is employed to modulate the electrical signal on an optical carrier at $\sim$1550 nm. The DC bias applied to the EAM is -1.6 V, and the insertion loss of the EAM is $\sim$9 dB.
After 2-km SSMF transmission, the received optical power (ROP) is adjusted by a variable optical attenuator (VOA). Then, the optical signal is converted to an electrical signal through a 30-Gbps PIN photodiode with an integrated trans-impedance amplifier (PIN-TIA). Finally, the electrical signal is converted to a digital signal by a 70 GSa/s analog-to-digital converter (ADC). The offline digital signal processing includes matched filter, resampling, synchronization, FFE, post-filter, MLSE, PAM4 demapping, and BER calculation. The estimated frequency domain channel response of this bandwidth-limited system is illustrated in the inset of Fig. 10. Under 2-km SSMF transmission, the 10-dB bandwidth of the whole system is only $\sim$15 GHz. Meanwhile, the frequency response drops rapidly beyond 21 GHz, which leads to severe bandwidth-limited distortion on the PAM4 signal.
Figure 10: Experiment setup of 112-Gbit/s PAM4 based on MLSE after 2-km SSMF transmission.
@@ VI-A Parameter Configuration for MLSE
Figure 11 illustrates different values of overlapped symbols $O$ for parallel MLSE. In our experiment, the number of received symbols $R$ is set to 16, since smaller values impair the performance of parallel MLSE. When the number of overlapped symbols is equal to or larger than 8, parallel MLSE can achieve almost the same BER performance as serial MLSE. When the number of overlapped symbols is less than 8, parallel MLSE has worse BER performance than serial MLSE. Therefore, in the parallel MLSE architecture of our paper, the number of symbols $N$ in a sliding block is set to 32, which includes 8 symbols of pre-overlap, 16 symbols of useful data, and 8 symbols of post-overlap. All the algorithms described in our paper are applied with this parallel architecture.
Figure 11: BER performance of parallel MLSE with different values of overlapped symbols $O$.
Figure 12: Experimental results of 112-Gbit/s PAM4 with FFE only and simplified L2S-MLSE after 2-km SSMF transmission.
@@ VI-B BER Performance, Latency and Complexity
After determining the parallel architecture, the experimental results of 112-Gbit/s PAM4 after 2-km SSMF transmission are shown in Fig. 12. As described in Section IV-B, the term simplified L2S-MLSE in our paper refers to L2S-MLSE with the combination of computational simplification and 2 reduced states. At the ROP of -7.5 dBm, the BER performance of 112-Gbit/s PAM4 after 2-km SSMF transmission achieves the 7% FEC limit with simplified L2S-MLSE, whereas it fails to meet the 7% FEC limit when only FFE is applied. If the ROP is further increased, the BER performance becomes worse due to the saturation of the TIA. With the symbol length $N=32$, the latency and complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE are shown in Fig. 13. L2S-MLSE architecture can reduce the latency from 34 delay units in linear order to 7 delay units in logarithmic order.
When considering computational complexity, both constant multipliers and comparators can be implemented using adders. Thus, variable multipliers occupy far more hardware resources than constant multipliers, adders, or comparators. In simplified 1S-MLSE and simplified L2S-MLSE, the number of variable multipliers is reduced from 512 in exponential order to 33 in linear order. Compared with L2S-MLSE, the number of adders and comparators for simplified L2S-MLSE is reduced to 37.2% and 8.4%, respectively. The number of adders and comparators in L2S-MLSE is 2.45 times and 3.88 times that of 1S-MLSE, respectively, while the number of adders and comparators in simplified L2S-MLSE drops to 1.80 times and 1.95 times that of simplified 1S-MLSE, respectively. In conclusion, the simplified scheme can reduce the utilization of multipliers, adders, and comparators.
Figure 13: Latency and computational complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE with $N=32$.
$BM_{n+1}(-3,-3)=\min\{A_{1}+B_{1},\hskip 7.0ptA_{2}+B_{2},\hskip 7.0ptA_{3}+B_{3},\hskip 7.0ptA_{4}+B_{4}\hskip 8.0pt\}$ (A.1)
$BM_{n+1}(-1,-3)=\min\{A_{5}+B_{5},\hskip 7.0ptA_{6}+B_{6},\hskip 7.0ptA_{7}+B_{7},\hskip 7.0ptA_{8}+B_{8}\hskip 8.0pt\}$ (A.2)
$BM_{n+1}(1,-3)\hskip 7.0pt=\min\{A_{9}+B_{9},\hskip 7.0ptA_{10}+B_{10},A_{11}+B_{11},A_{12}+B_{12}\hskip 1.0pt\}$ (A.3)
$BM_{n+1}(3,-3)\hskip 7.0pt=\min\{A_{13}+B_{13},A_{14}+B_{14},A_{15}+B_{15},A_{16}+B_{16}\hskip 1.0pt\}$ (A.4)
$BM_{n+1}(-3,-1)=\min\{C_{1}+D_{1},\hskip 7.0ptC_{2}+D_{2},\hskip 7.0ptC_{3}+D_{3},\hskip 7.0ptC_{4}+D_{4}\hskip 8.0pt\}$ (A.5)
$BM_{n+1}(-1,-1)=\min\{C_{5}+D_{5},\hskip 7.0ptC_{6}+D_{6},\hskip 7.0ptC_{7}+D_{7},\hskip 7.0ptC_{8}+D_{8}\hskip 8.0pt\}$ (A.6)
$BM_{n+1}(1,-1)\hskip 7.0pt=\min\{C_{9}+D_{9},\hskip 7.0ptC_{10}+D_{10},C_{11}+D_{11},C_{12}+D_{12}\}$ (A.7)
$BM_{n+1}(3,-1)\hskip 7.0pt=\min\{C_{13}+D_{13},C_{14}+D_{14},C_{15}+D_{15},C_{16}+D_{16}\}$ (A.8)
$BM_{n+1}(-3,1)\hskip 7.0pt=\min\{C_{16}-D_{16},C_{15}-D_{15},C_{14}-D_{14},C_{13}-D_{13}\}$ (A.9)
$BM_{n+1}(-1,1)\hskip 7.0pt=\min\{C_{12}-D_{12},C_{11}-D_{11},C_{10}-D_{10},C_{9}-D_{9}\hskip 7.0pt\}$ (A.10)
$BM_{n+1}(1,1)\hskip 14.0pt=\min\{C_{8}-D_{8},\hskip 8.0ptC_{7}-D_{7},\hskip 7.0ptC_{6}-D_{6},\hskip 7.0ptC_{5}-D_{5}\hskip 7.0pt\}$ (A.11)
$BM_{n+1}(3,1)\hskip 14.0pt=\min\{C_{4}-D_{4},\hskip 8.0ptC_{3}-D_{3},\hskip 7.0ptC_{2}-D_{2},\hskip 7.0ptC_{1}-D_{1}\hskip 7.0pt\}$ (A.12)
$BM_{n+1}(-3,3)\hskip 7.0pt=\min\{A_{16}-B_{16},A_{15}-B_{15},A_{14}-B_{14},A_{13}-B_{13}\hskip 1.0pt\}$ (A.13)
$BM_{n+1}(-1,3)\hskip 7.0pt=\min\{A_{12}-B_{12},A_{11}-B_{11},A_{10}-B_{10},A_{9}-B_{9}\hskip 8.0pt\}$ (A.14)
$BM_{n+1}(1,3)\hskip 14.0pt=\min\{A_{8}-B_{8},\hskip 8.0ptA_{7}-B_{7},\hskip 7.0ptA_{6}-B_{6},\hskip 7.0ptA_{5}-B_{5}\hskip 8.0pt\}$ (A.15)
$BM_{n+1}(3,3)\hskip 14.0pt=\min\{A_{4}-B_{4},\hskip 8.0ptA_{3}-B_{3},\hskip 7.0ptA_{2}-B_{2},\hskip 7.0ptA_{1}-B_{1}\hskip 8.0pt\}$ (A.16)

### v2

@@ IV Simplified MLSE
Since the latency of MLSE is reduced at the cost of complexity, simplified MLSE is proposed in this section. Simplified MLSE consists of two parts: computational simplification and reduced state. Computational simplification is proposed to reduce the number of multipliers. Reduced state is combined with computational simplification, further reducing the number of adders and comparators.
@@ IV-A Computational Simplification
1) BM Calculation and ACS of the First Layer: To reduce the number of multipliers, the computational simplification is proposed, which combines the common terms in the BM calculation and addition operations of the first layer. The 64 expansion equations of the BM calculation and addition operations from time index n1n-1 to time index $n+1$ in the first layer are shown in Appendix A. These 64 expansion equations can be expressed as the sum of vectors $\bf{A}$}, $\bf{B}$}, $\bf{C}$}, and $\bf{D}$}. Both the vector ${\bf{A}}=[{A_{1}},{A_{2}},...,{A_{16}}]$ and the vector ${\bf{C}}=[{C_{1}},{C_{2}},...,{C_{16}}]$ are related to $\alpha$ and constant values. Both the vector ${\bf{B}}=[{B_{1}},{B_{2}},...,{B_{16}}]$ and the vector ${\bf{D}}=[{D_{1}},{D_{2}},...,{D_{16}}]$ are related to the received symbol $y$.
@@
Through combining the common terms, the vectors $\bf{A}$} and $\bf{C}$} include 1 variable multiplier for $\alpha^{2}$, 4 constant multipliers for $10\alpha^{2}$, $18\alpha^{2}$, $12\alpha$, $36\alpha$, and additional 31 adders. Multiplication with powers of 2 can be achieved through bit-shift operations, thereby avoiding the use of multipliers. For the calculation of $\bf{B}$} and $\bf{D}$}, $\bf{B}$} can be decomposed into ${\bf{M}}+{\bf{N}}+{\bf{J}}$, while $\bf{D}$} can be decomposed into ${\bf{M}}+{\bf{N}}+{\bf{K}}$. $\bf{M}$} corresponds to the common term $y_{n}+\alpha y_{n+1}$ and its multiples. $\bf{N}$} corresponds to the common term $\alpha y_{n}$ and its multiples. $\bf{J}$} is equal to the common term $6y_{n+1}$ for $\bf{B}$}, and $\bf{K}$} is equal to the common term $2y_{n+1}$ for $\bf{D}$}.
@@
Therefore, through combining the common terms, the vectors $\bf{B}$} and $\bf{D}$} include 2 variable multipliers, 3 constant multipliers, and 41 adders. The computation of $\bf{B}$} and $\bf{D}$} can be performed simultaneously with that of $\bf{A}$} and $\bf{C}$}. In conclusion, with computational simplification, the overall complexity of BM calculation and ACS operation in the first layer includes $N+1$ variable multipliers, $3N/2+4$ constant multipliers, $105N/2+31$ adders, and $24N$ comparators.
@@
2) ACS of the Other Layers and Survivor Path Selection: The difference between MLSE with and without computational simplification only exists in the first layer, since the second layer can only utilize the computationally simplified results from the first layer. From the second layer to the last layer, the computing process of MLSE with computational simplification is identical to that of MLSE without computational simplification, which preserves the same latency and computational complexity.
@@
To make a fair comparison, the computational simplification is also applied to 1S-MLSE. The expansion equations can be expressed as the sum of vectors $\bf{E}$} and $\bf{F}$}. By combining common terms, the vector ${\bf{E}}=[{E_{1}},{E_{2}},...,{E_{8}}]$ includes 1 variable multiplier for $\alpha^{2}$, 3 constant multipliers for $9\alpha^{2}$, $6\alpha$, $18\alpha$, and 12 adders. The vector ${\bf{F}}=[{F_{1}},{F_{2}},...,{F_{8}}]$ includes 1 variable multiplier for $\alpha y_{n}$, 2 constant multipliers for $6\alpha y_{n}$ and $6y_{n}$, plus 8 adders. 16 adders are needed for the addition between $\bf{E}$} and $\bf{F}$}, while 12 comparators are utilized after the addition operations. After the BM calculation with computational simplification, ACS and survivor path selection are the same as those of 1S-MLSE without computational simplification, which preserves the same latency and computational complexity.
@@
In conclusion, the overall computational complexity of 1S-MLSE with computational simplification is $N+1$ variable multipliers, $2N+3$ constant multipliers, 40N440N-4 adders, and $12N+3$ comparators, while that of MLSE with computational simplification is $N+1$ variable multipliers, $3N/2+4$ constant multipliers, 169N/233169N/2-33 adders, and 48N3348N-33 comparators. Both 1S-MLSE with computational simplification and MLSE with computational simplification use the same number of variable multipliers. Because constant multipliers and comparators can be implemented by adders, variable multipliers occupy far more hardware resources than constant multipliers, adders, or comparators. When using computational simplification, the number of variable multipliers is reduced from $O(NM^{L})$ to $O(N)$. The BER performance of 1S-MLSE, 1S-MLSE with computational simplification, MLSE, and MLSE with computational simplification is analyzed. All four algorithms have the same BER performance. By combining common items, the calculation accuracy remains unchanged, ensuring the computational simplification does not degrade the BER performance.
@@
IV-B Simplified MLSE with the Combination of Computational Simplification and Reduced State
@@
To reduce the number of adders and comparators, the reduced state is combined with computational simplification. In our paper, we refer to MLSE with the combination of computational simplification and reduced state as simplified MLSE. The output signal $d$ of FFE can be regarded as the pre-decision value of the received signal [26]. For PAM4 with 4 states of $\{-3,-1,1,3\}$, if the number of states is reduced from 4 to 3 through pre-decision value, the possible states include $\{-3,-1,1\}$ or $\{-1,1,3\}$, which are represented by $\{a,b,c\}$. If the number of states is reduced from 4 to 2 through pre-decision value, the possible states include $\{-3,-1\}$, $\{-1,1\}$ or $\{1,3\}$, which are represented by $\{a,b\}$. When the state number is reduced to 3, the number of ACS units in a calculation is reduced from 16 to 9, in which only 27 adders and 18 comparators are needed. When the state number is reduced to 2, the number of ACS units in a calculation is reduced from 16 to 4, in which only 8 adders and 4 comparators are needed.
@@
2) ACS of the Other Layers and Survivor Path Selection: From the second layer to the last layer, due to the decrease in ACS units, 27 adders and 18 comparators are needed for a calculation with 3 states. 8 adders and 4 comparators are needed for a calculation with 2 states. In the final survivor path selection step, the path with the minimal accumulated metrics is chosen from 9 paths for the state number of 3, reducing the number of comparators to 8. Similarly, the path with the minimal accumulated metrics is chosen from 4 paths for the state number of 2, reducing the number of comparators to 3.
We compares the BER performance of MLSE under three configurations: all 4 states (with only computational simplification), simplified MLSE with 3 states, and simplified MLSE with 2 states. In our experiment, the noise enhancement caused by FFE is not severe enough to damage the pre-decision value from FFE for 3 or 2 reduced states. Therefore, the BER performance remains nearly identical regardless of whether all 4 states, 3 states, or 2 states are used. For the system with a larger $\alpha$ value, due to the larger noise in the pre-decision value, the BER performance of 2 states may have worse BER performance than 3 states [26]. Owing to the nearly identical BER performance, the simplified scheme in our paper refers to the combination of computational simplification and 2 reduced states. To make a fair comparison, the simplified scheme is also applied to 1S-MLSE.
@@ V Experimental Setup
The experiment of a 112-Gbit/s PAM4 signal over 2-km SSMF transmission in C-band is performed to verify the performance of the proposed MLSE. At the transmitter, $4\times 10^{5}$ pseudo-random binary sequences (PRBS) are mapped to $2\times 10^{5}$ PAM4 symbols, in which the first 1000 symbols are served as training symbols. To improve the transmission performance, the optimized pre-equalization method proposed in our previous work is applied to compensate for partial bandwidth-limited distortion [10]. The PAM4 signal is pulse-shaped by a root-raised cosine (RRC) filter with the roll-off factor of 0.1. Then, the signal is resampled by 1.25 sps to align with the sampling rate of the digital-to-analog converter (DAC). The 56-GBaud PAM4 analog signal is generated by a DAC with a sampling rate of 70 GSa/s. The 3-dB bandwidth of the 8-bit DAC is 16 GHz. Afterwards, the analog signal is amplified by a 40-Gbps electrical amplifier (EA, CENTELLAX OA4MVM3). A 40-Gbps optical electro-absorption modulator (EAM, OM5757C-CTM388) is employed to modulate the electrical signal on an optical carrier at $\sim$1550 nm. The DC bias applied to the EAM is -1.6 V, and the insertion loss of the EAM is $\sim$9 dB.
After 2-km SSMF transmission, the received optical power (ROP) is adjusted by a variable optical attenuator (VOA). Then, the optical signal is converted to an electrical signal through a 30-Gbps PIN photodiode with an integrated trans-impedance amplifier (PIN-TIA). Finally, the electrical signal is converted to a digital signal by a 70 GSa/s analog-to-digital converter (ADC). The offline digital signal processing includes matched filter, resampling, synchronization, FFE, post-filter, MLSE, PAM4 demapping, and BER calculation. Under 2-km SSMF transmission, the 10-dB bandwidth of the whole system is only $\sim$15 GHz. Meanwhile, the frequency response drops rapidly beyond 21 GHz, which leads to severe bandwidth-limited distortion on the PAM4 signal.
@@ VI-A Parameter Configuration for MLSE
In our experiment, the number of received symbols $R$ is set to 16, since smaller values impair the performance of parallel MLSE. When the number of overlapped symbols is equal to or larger than 8, parallel MLSE can achieve almost the same BER performance as serial MLSE. When the number of overlapped symbols is less than 8, parallel MLSE has worse BER performance than serial MLSE. Therefore, in the parallel MLSE architecture of our paper, the number of symbols $N$ in a sliding block is set to 32, which includes 8 symbols of pre-overlap, 16 symbols of useful data, and 8 symbols of post-overlap. All the algorithms described in our paper are applied with this parallel architecture.
@@ VI-B BER Performance, Latency and Complexity
After determining the parallel architecture, the experimental results of 112-Gbit/s PAM4 after 2-km SSMF transmission are analyzed. As described in Section IV-B, the term simplified MLSE in our paper refers to MLSE with the combination of computational simplification and 2 reduced states. At the ROP of -7.5 dBm, the BER performance of 112-Gbit/s PAM4 after 2-km SSMF transmission achieves the 7% FEC limit with simplified MLSE, whereas it fails to meet the 7% FEC limit when only FFE is applied. If the ROP is further increased, the BER performance becomes worse due to the saturation of the TIA. With the symbol length $N=32$, the latency and complexity of 1S-MLSE, simplified 1S-MLSE, MLSE, and simplified MLSE are analyzed. MLSE architecture can reduce the latency from 34 delay units in linear order to 7 delay units in logarithmic order.
When considering computational complexity, both constant multipliers and comparators can be implemented using adders. Thus, variable multipliers occupy far more hardware resources than constant multipliers, adders, or comparators. In simplified 1S-MLSE and simplified MLSE, the number of variable multipliers is reduced from 512 in exponential order to 33 in linear order. Compared with MLSE, the number of adders and comparators for simplified MLSE is reduced to 37.2% and 8.4%, respectively. The number of adders and comparators in MLSE is 2.45 times and 3.88 times that of 1S-MLSE, respectively, while the number of adders and comparators in simplified MLSE drops to 1.80 times and 1.95 times that of simplified 1S-MLSE, respectively. In conclusion, the simplified scheme can reduce the utilization of multipliers, adders, and comparators.